The supply function of a good is given by Ps=Q2-2Q+12. Determine the producer surplus if ten units are supplied (Q=10)at price P.


1
Expert's answer
2021-07-14T12:02:14-0400

If Q = 10, then "P = 10^2 - 2*10 + 12 = 92."

If Q = 0, then P = 12.

The producer surplus is: "PS = 0.5*10*(92 - 12) = 400."